out of 300 families with 5 children each, what percentage would be expected to have

i. at least a boy

ii. all boys

iii. at most 2 girls


1
Expert's answer
2021-06-21T10:11:47-0400

Let binomial random variable "X" denotes the number of boys in the family: "X\sim B(n, p)."

Since boy and girl child have equal probability, then "p=q=0.5."

Given "n=5." Then "X\sim B(5, 0.5)"


i.

"P(X\geq1)=1-P(X<1)=1-P(X=0)"

"=1-\dbinom{5}{0}(0.5)^0(0.5)^{5-0}=0.96875"

"96.875\ \%"


ii.

"P(X=5)=\dbinom{5}{0}(0.5)^5(0.5)^{5-5}=0.03125"

"3.125\ \%"


iii.

"P(X\geq3)=P(X=3)+P(X=4)+P(X=5)"

"=\dbinom{5}{3}(0.5)^3(0.5)^{5-3}+\dbinom{5}{4}(0.5)^4(0.5)^{5-4}"




"+\dbinom{5}{5}(0.5)^5(0.5)^{5-5}=(10+5+1)(0.5)^5=0.5"

"50\ \%"